OpenStudy (anonymous):

The sum of a number and two times a smaller number is 62. Three times the bigger number exceeds the smaller number by 116. The bigger number is . The smaller number is .

OpenStudy (campbell_st):

let x = big number and y = small number create 2 equations x + 2y = 62 3x - y = 116 now you can solve them silumtaneously

OpenStudy (anonymous):

bigger number: 42 smaller number: 10 use systems of equations: y+2x=62 3y=x+116